You are selling tickets for a high school basketball game. Student
tickets cost $3 and general admission tickets cost $5. You sell 350
tickets and collect $1,450. How many general admission tickets did
you sell?

Answer by checkley79(3341)

5x+3(350-x)=1,450
5x+1,050-3x=1,450
5x-3x=1,450-1,050
2x=400
x=400/2
x=200 $5.00 general admission tickets were sold.
350-200=150 $ 3.00 tickets were sold.
Proof:
5*200+3*150=1,450
1,000+450=1,450
1,450=1,450